Operations Report – May 24th

Crew 299 Operations Report 5-24-2024

SOL: 12

Name of person filing report: Rishabh Pandey (Engineer)

Non-Nominal Systems: Suit 9 (broken visor), Suit 11 (fan issue)

Notes on non-nominal systems: Issues with both suits were noted before our mission, both will not be used during our mission.

Rovers Used: None
General notes on rovers: None

Summary of Hab operations:
      WATER USE: 41 Gallons
      Water (static tank): 178.3 (using effective value from spreadsheet) (was also refilled today)
      Static tank pipe heater: off
      Static tank heater: off
      Toilet tank emptied: yes

Summary of internet usage: Catch up on Earthly news, email responses, code development, entertainment

Summary of suits and radios: No suits or radios used.

Summary of GreenHab operations: NA, GreenHab out of operation

Summary of ScienceDome operations: Final soil analysis was conducted.

Summary of RAM operations: Measuring tape used.
Summary of observatory issues: NA

Copyright © The Mars Society. All rights reserved. | Main Site